# Eaman Jahani

### v3

@@ Abstract
The recent surge in generative AI has led to new models being introduced almost every month. In light of this rapid progression, we pose and address a central question: to what extent must prompts evolve as the capabilities of generative AI models advance? To answer this question, we conducted an online experiment with N = 1,893 participants where each participant was incentivized to write prompts to reproduce a target image as closely as possible in 10 consecutive tries. Each participant was randomly and blindly assigned to use one of three text-to-image diffusion models: DALL-E 2, its more advanced successor, DALL-E 3, or a version of DALL-E 3 with automatic prompt revision. In total, we collected and analyzed over 18,000 prompts and over 300,000 images. We find that task performance was higher for participants using DALL-E 3 than for those using DALL-E 2. This performance gap corresponds to a noticeable difference in the similarity of participants’ images to their target images, and was caused in equal measure by: (1) the increased technical capabilities of DALL-E 3, and (2) endogenous changes in participants’ prompting in response to these increased capabilities. Furthermore, while participants assigned to DALL-E 3 with prompt revision still outperformed those assigned to DALL-E 2, automatic prompt revision reduced the benefits of using DALL-E 3 by 58%. Our results suggest that for generative AI to realize its full impact on the global economy, people, firms, and institutions will need to update their prompts in response to new models. Not doing so could leave more than half of the potential benefits of these AI systems untapped.
@@
Recent economic forecasts predict that generative AI could add trillions of dollars to the global economy annually, and there is mounting evidence that this new technology is being integrated into work practices in areas as diverse as business, medicine, and government FactSet (2024); Zhang and Kamel Boulos (2023); Bright et al. (2024). This increasing generative AI adoption appears to be warranted, with recent studies finding that the adoption of generative AI improves the productivity of workers across various professional settings (Brynjolfsson et al., 2023; Dell’Acqua et al., 2023; Noy and Zhang, 2023). A common feature of these studies is that the model participants interacted with did not change; however, generative models are continuously updated, with new versions being released as often as once a month. Humans interface with generative models by providing textual instructions or “prompts” to the models (Oppenlaender, 2023; Schulhoff et al., 2024; Don-Yehiya et al., 2023; Xie et al., 2023). It remains an open question whether organizations will need to adapt these prompts as models improve to realize the full economic and societal benefits of generative AI. For example, it is now common amongst software companies to have a hundred or more development teams, each with dozens, if not hundreds, of prompts integrated with their code. On the one hand, firms might write these prompts once and subsequently benefit from new models with minimal additional effort. Alternatively, as models improve, organizations may need to continually adapt their prompts. In this paper, we investigate whether this prompt adaptation process is necessary, and the extent to which the ability to adapt one’s prompts is a specialized skill.
@@
We do so by conducting a pre-registered online lab experiment in which $N=1,893{}$ participants from Prolific were randomly and blindly assigned to complete the same task using one of three generative AI models with varying capabilities: (1) DALL-E 2, (2) its more advanced successor DALL-E 3, or (3) a version of DALL-E 3 with automatic large language model (LLM)-based prompt revisions (hereafter referred to as “DALL-E 3 with revision”). [^1] The main text presents a subset of pre-registered analyses, with additional results and deviations from pre-registration detailed in the SI. The LLM-based prompt revisions in the third condition were not made visible to participants during the experiment.222prompt revision is the default behavior of the DALL-E 3 API endpoint. We followed the DALL-E 3 API documentation and attempted to disable prompt revision in the second, “DALL-E 3” treatment arm by prepending a system prompt to participants’ prompts. While this drastically reduced the number and extent of LLM-based prompt revisions, it did not completely eliminate them. Therefore, the “DALL-E 3” treatment arm still includes some automatic prompt revisions. For brevity and simplicity, we refer to this treatment arm as “DALL-E 3” throughout the paper. Full details on the system prompt’s efficacy are provided in the SI. The task required each participant to make at least 10 attempts at recreating a “target image” as closely as possible by prompting their assigned model. Each participant’s target image was randomly selected from a curated set of 15 images; this set of images was constructed to span realistic use cases of AI-generated images. Throughout the task, participants could view all of their past prompts and generated images alongside the target image. Participants were paid \$4 USD to complete the task, and to incentivize performance, they received \$8 USD if they were in the top 20% of participants, as measured by the similarity of their best attempt to the target image.
@@ Results
The changes in prompting we observe are an important mechanism through which participant performance improves, suggesting that organizations will need to adapt their prompts as models evolve. More specifically, using the decomposition procedure described in the Materials and Methods, we find that changes in prompting accounted for 48% of the performance ATE ($\Delta CoSim=0.00788$, $p=0.024$; blue in Figure 2), whereas the shift to a more capable model accounted for 51% of the performance ATE ($\Delta CoSim=0.00841$, $p<10^{-8}$; red in Figure 2).
We also supplied the prompts written for DALL-E 3 to DALL-E 2 (see Materials and Methods for details). Doing so did not improve performance ($\Delta CoSim=0.0020$, $p=0.56$), indicating that DALL-E 3 prompts specifically took advantage of its superior ability to render information (Betker et al., 2023).
@@
Our results show that adapting prompts is crucial for maximizing the benefits of more capable generative models, raising the question of how organizations should approach this adaptation process. One possibility is that LLM-based, automated prompt revision (Li et al., 2024; Betker et al., 2023) obviates the need for humans to tailor their prompts to different generative models. Comparison of our third treatment arm–DALL-E 3 with revision–to the DALL-E 3 arm highlights the potential pitfalls of this approach. We find that prompt revision actually caused participants to perform worse in our context as it reduced the benefit of using DALL-E 3 by nearly 58% (95% CI: [40%, 76%]).
Furthermore, the positive impact of access to DALL-E 3 with revision performed only slightly better than those assigned to DALL-E 2 ($\Delta CoSim=0.0069$; $p=0.042$) and was less than the positive impact of directly passing prompts written for DALL-E 2 to DALL-E 3 (i.e., the model effect shown in Figure 2). These findings suggest that as currently implemented, AI-assisted prompt revisions are not a panacea. They can actually inhibit people’s capacity to leverage a model’s capabilities when misaligned with an end user’s goals.
@@ Discussion
The primary limitation of our work is that we only studied the transition from DALL-E 2 to DALL-E 3. We leave the study of future transitions and different types of generative AI models to future work. Should our results generalize to other models, then we would predict that firms that do not update their prompts as models progress are leaving roughly half the benefit of AI unrealized. Software systems, models, and prompts are becoming increasingly interwoven which is a trend we expect to accelerate. Thus, the lack of prompt maintenance could, in turn, dampen the economic impact of AI on the global economy. Fortunately, updating prompts can be learned by humans across the skill distribution even if models are not currently able to update prompts in an automated fashion. Thus, we anticipate that as firms adopt generative AI, they will need to regularly refine prompts to fully realize the potential of evolving models. In particular, one could imagine a lock-step dynamic where, as models continually improve, people will need to—and are easily able to—respond by adapting their prompts to take advantage of the newest model’s capabilities. Such a pattern suggests that as generative AI models advance, technical prompting infrastructure will not be a one-time investment. Rather, prompting will be the mechanism by which people and firms unlock new models’ capabilities.
Table 1: Model and Prompting Effect by Performance
Effect Estimate p-value
Model $0.011000$ $p<0.00001$
$(0.001860)$
Model $\times$ $-0.000060$ $0.0210$
Performance Decile $(0.000030)$
Prompting $0.010900$ $0.0006$
$(0.003160)$
Prompting $\times$ $-0.000060$ $0.2444$
Performance Decile $(0.000050)$
@@
Outcome measure Given the stochastic nature of generative AI models, our outcome measure of task performance is the expected similarity of images generated from each prompt to the corresponding target image. To compute this, for each prompt we generated 10 images and computed CLIP embeddings (Radford et al., 2021) for both these and the target images. Each prompt’s expected similarity score was then computed as the mean cosine similarity ($CoSim$) between its generated image embeddings and the target image embedding (see SI for details).
Effect Decomposition We conducted an exploratory analysis where we regenerated images with all participant prompts on both models. This allows us to decompose the ATE into two components: the “model effect” (the average improvement when running DALL-E 2 prompts through DALL-E 3) and the “prompting effect” (the difference in target similarity between DALL-E 3 and DALL-E 2 participants’ prompts, both evaluated on DALL-E 3). Note that all main text analyses use “replay” data from Figure 2 for consistency and to avoid model drift issues. All results hold with the “original” data. See the SI for details.
Experimental Materials Preregistration, data, and analysis code will be deposited at https://osf.io/ejbtp. Full details regarding our experiment design, analysis techniques, and additional results can be found in the SI.

### v4

@@ Abstract
The recent surge in generative AI has led to new models being introduced almost every month. In light of this rapid progression, we pose and address a central question: to what extent must prompts evolve as the capabilities of generative AI models advance? To answer this question, we conducted an online experiment with N = 1,893 participants where each participant was incentivized to write prompts to reproduce a target image as closely as possible in 10 consecutive tries. Each participant was randomly and blindly assigned to use one of three text-to-image diffusion models: DALL-E 2, its more advanced successor, DALL-E 3, or a version of DALL-E 3 with automatic prompt revision. In total, we collected and analyzed over 18,000 prompts and over 300,000 images. We find that task performance was higher for participants using DALL-E 3 than for those using DALL-E 2. This performance gap corresponds to a noticeable difference in the similarity of participants’ images to their target images, and was caused in equal measure by: (1) the increased technical capabilities of DALL-E 3, and (2) endogenous changes in participants’ prompting in response to these increased capabilities. Furthermore, while participants assigned to DALL-E 3 with prompt revision still outperformed those assigned to DALL-E 2, automatic prompt revision reduced the benefits of using DALL-E 3 by 58%. Our results suggest that for generative AI to realize its full impact on the global economy, people, firms, and institutions will need to update their prompts in response to new models.
@@
Recent economic forecasts predict that generative AI could add trillions of dollars to the global economy annually, and there is mounting evidence that this new technology is being integrated into work practices in areas as diverse as business, medicine, and government FactSet (2024); Zhang and Kamel Boulos (2023); Bright et al. (2024). This widespread adoption appears to be warranted, with recent studies finding that giving workers generative AI improves productivity across various professional settings (Brynjolfsson et al., 2023; Dell’Acqua et al., 2023; Noy and Zhang, 2023). A common feature of these studies is that they used static model versions for testing. However, in practice, generative models are continuously updated, with new versions being released as frequently as once per month. People interact with generative AI by providing textual instructions or “prompts” to guide the models’ outputs (Oppenlaender, 2023; Schulhoff et al., 2024; Don-Yehiya et al., 2023; Xie et al., 2023). It remains unclear whether organizations must continually adjust these prompts as the models advance to fully capture their economic and societal advantages. For example, it is now common amongst software companies to have a hundred or more development teams, each with hundreds, if not thousands, of prompts integrated with their code. On the one hand, firms might write these prompts once and subsequently benefit from new models with minimal additional effort. Alternatively, organizations may need to continually adapt their prompts as models improve. In this paper, we investigate whether refining prompts is necessary as models evolve, and how much specialized expertise this refining process requires.
@@
To explore these questions, we conducted a pre-registered online lab experiment. $N=1,893{}$ participants from Prolific were randomly and blindly assigned to complete the same task using one of three generative AI models with varying capabilities: (1) DALL-E 2, (2) its more advanced successor DALL-E 3, or (3) a version of DALL-E 3 with automatic large language model (LLM)-based prompt revisions (hereafter referred to as “DALL-E 3 with revision”). [^1] The main text presents a subset of pre-registered analyses, with additional results and deviations from pre-registration in the SI. The LLM-based prompt revisions in the third condition were not visible to participants during the experiment. [^2] Automatic prompt revision is the default behavior of the DALL-E 3 API. These prompt revisions involved sending each participant’s prompt to GPT-4, which enhanced it by adding descriptive details before being sent to DALL-E 3. We followed the DALL-E 3 API documentation and attempted to disable prompt revision in the second, “DALL-E 3” treatment arm by prepending a system prompt to participants’ prompts. While this drastically reduced the number and extent of LLM-based prompt revisions, it did not completely eliminate them. Therefore, the “DALL-E 3” treatment arm still includes some automatic revisions. We refer to this treatment arm as “DALL-E 3” for brevity and simplicity throughout the paper. Full details on the system prompt’s efficacy are provided in the SI. The task required each participant to make at least 10 attempts at recreating a “target image” as closely as possible by prompting their assigned model. Each participant’s target image was randomly selected from a curated set of 15 images; this set of images was constructed to span realistic use cases of AI-generated images. Throughout the task, participants could view all of their past prompts and generated images alongside the target image. Participants were paid \$4 USD to complete the task, and to incentivize performance, they received \$8 USD if they were in the top 20% of participants, as measured by the similarity of their best attempt to the target image.
@@ Results
The changes in prompting we observe are an important mechanism through which participant performance improves, suggesting that organizations must adapt their prompts as models evolve. To quantify how much of the performance improvement came from participants adapting their prompts, we took prompts written by participants in the DALL-E 2 treatment arm and generated images with them using DALL-E 3. This allowed us to isolate the “model effect”—how much better DALL-E 3 performed with prompts written for DALL-E 2 ($\Delta CoSim=0.0084$, $p<10^{-8}$; red in Figure 2), and the “prompting effect”—the portion of the performance ATE due solely to participants adapting their prompting for DALL-E 3 ($\Delta CoSim=0.0079$, $p=0.024$; blue in Figure 2). These effects were nearly equal: 51% and 48% of the performance ATE, respectively. In contrast, we also supplied the prompts written for DALL-E 3 to DALL-E 2. Doing so did not improve performance ($\Delta CoSim=0.0020$, $p=0.56$). This implies that DALL-E 3 prompts specifically took advantage of its superior ability to render information, and users naturally discovered these capabilities (Betker et al., 2023).
@@
Our results show that adapting prompts is crucial for maximizing the benefits of more capable generative models, raising the question of how organizations should approach this adaptation process. One possibility is that LLM-based, automated prompt revision (Li et al., 2024; Betker et al., 2023) obviates the need for humans to tailor their prompts to different generative models. Comparison of our third treatment arm—DALL-E 3 with revision—to the DALL-E 3 arm highlights the potential pitfalls of this approach. We find that automated prompt revisions actually caused participants to perform worse in our context. It reduced the benefit of using DALL-E 3 by nearly 58% (95% CI: [40%, 76%]). Furthermore, the positive impact of access to DALL-E 3 with revision performed only slightly better than those assigned to DALL-E 2 ($\Delta CoSim=0.0069$; $p=0.042$) and was less than the positive impact of directly passing prompts written for DALL-E 2 to DALL-E 3 (i.e., the model effect shown in Figure 2). These findings suggest that, as currently implemented, AI-assisted prompt revisions are not a panacea. They can actually inhibit people’s capacity to leverage a model’s capabilities when misaligned with an end user’s goals. Humans, on the other hand, are better at updating their prompts.
@@
Table 1: Model and Prompting Effect by Performance Decile
Effect Estimate P-value
Model $0.01100$ $<0.00001$
$(0.00186)$
Model $\times$ $-0.00006$ $0.0210$
Performance Decile $(0.00003)$
Prompting $0.01090$ $0.0006$
$(0.00316)$
Prompting $\times$ $-0.00006$ $0.2444$
Performance Decile $(0.00005)$
@@ Discussion
The main limitation of our work is that we only study the transition from DALL-E 2 to DALL-E 3. We leave the study of future transitions and different types of generative AI models to future research. Software systems, models, and prompts are becoming increasingly interwoven, which is a trend we expect to accelerate. If our findings generalize to other models, firms that fail to update their prompts as models improve may forego roughly half of the potential benefits of AI advancement, which will dampen the overall economic impact of AI. Thus, we anticipate that as firms adopt generative AI, they will need to regularly refine prompts to fully realize the potential of evolving models. In particular, one could imagine a lock-step dynamic where, as models continually improve, people will need to—and are easily able to—respond by adapting their prompts to take advantage of the newest model’s capabilities. Such a pattern suggests that technical prompting infrastructure will not be a one-time investment as generative AI models advance. Rather, prompting will be an essential and increasingly intuitive mechanism by which people and firms unlock new models’ capabilities. Fortunately, our findings suggest that humans across the skill distribution can learn to update prompts, even when automated revisions do not suffice.
@@
Institutional Review. This study was reviewed by the UC Berkeley Committee for the Protection of Human Subjects (CPHS) under Protocol 2023-06-16480.
Outcome Measure. Given the stochastic nature of generative AI models, our outcome measure of task performance is the expected similarity of images generated from each prompt to the corresponding target image. To compute this, we generated 10 images for each prompt and computed CLIP embeddings (Radford et al., 2021) for both these and the target images. Each prompt’s expected similarity score was then computed as the mean cosine similarity ($CoSim$) between its generated image embeddings and the target image embedding (see SI for details).
Replay Data. To generate the images used to decompose the prompting and model effects, we regenerated or “replayed” all images with all participant prompts on both models. Note that all main text analyses use replay data from Figure 2—even replays of DALL-E 2 written prompts on DALL-E 2, and DALL-E 3 written prompts on DALL-E 3—for consistency and to avoid model drift issues. All results hold with the “original” data. See SI for details.
Experimental Materials. Preregistration, data, and analysis code will be deposited at https://osf.io/ejbtp. Full details regarding our experiment design, analysis techniques, and additional results can be found in the SI.